OpenStudy (howard-wolowitz):

State the various transformation applied to the base function ........... to obtain a graph of the function ............

OpenStudy (howard-wolowitz):

\[f(x) = x^{3}]

OpenStudy (howard-wolowitz):

\[g(x) = 3(x+4)^{3} -5\]

OpenStudy (anonymous):

this might help http://www.purplemath.com/modules/fcntrans.htm

OpenStudy (howard-wolowitz):

Is this clear?

OpenStudy (michele_laino):

yes!

OpenStudy (howard-wolowitz):

A. Horizontal shift of 4 units to the left, a vertical stretch by a factor of 3, and a vertical shift downward of 15 units. B. Horizontal shift of 4 units to the left, a vertical stretch by a factor of 3, and a vertical shift downward of 5 units. C. Horizontal shift of 12 units to the left, a vertical stretch by a factor of 3, and a vertical shift downward of 5 units. D. Horizontal shift of 4 units to the right, a vertical stretch by a factor of 3, and a vertical shift downward of 5 units.

OpenStudy (michele_laino):

Please note that if I apply these traslations: \[\begin{gathered} Y = y + 5 \hfill \\ X = x + 4 \hfill \\ \end{gathered} \] where y=g(x), then I can write: \[Y = 3{X^3}\]

OpenStudy (michele_laino):

so I start from the function g(x), and I perform the above traslation, and finally I get the function f(x)

OpenStudy (howard-wolowitz):

umm so its not C.

OpenStudy (howard-wolowitz):

So I"m thinking B?

OpenStudy (michele_laino):

Yes! I think B